OpenStudy (anonymous):

Find The Domain f(x)=log(15-x) ANSWER IN INTERVAL NOTATION (x,y) Find The Domain f(x)=log(15-x) ANSWER IN INTERVAL NOTATION (x,y) @Mathematics

OpenStudy (anonymous):

guess the last answer didn't make sense. x has to be less than 15.

OpenStudy (anonymous):

i think the answer would be: \[x \in (-\infty, 0)\]

OpenStudy (anonymous):

it would be if it was the domain of \[\ln(-x)\] but since it is the domain of \[\ln(15-x)\] it is not \[(-\infty,0)\]

OpenStudy (anonymous):

you're right, silly mistake, so it would be \[x \in(-\infty, 15)\] i hope i got it right this time